# Unveiling the Role of ChatGPT in Software Development: Insights from Developer-ChatGPT Interactions on GitHub

### v2

@@ 6. Threats to Validity
In this section, we discuss the potential threats to the validity related to our study and the adopted measures mitigating these threats according to the guidelines of Wohlin et al. (Wohlin et al., 2012). Internal validity is not considered in this work, as no causal relationships between variables are addressed.
Construct validity: Potential threats of this study lie in the mined ChatGPT links and their content that developers shared on GitHub. Not all shared ChatGPT links reflect meaningful interactions, and some may be noise, off-topic, or duplicates. For example, many shared ChatGPT links are not relevant to software development, or certain development-related ChatGPT links are duplicated from different sources (e.g., one link occurs in both Code and Issues). To mitigate these threats, we employed rigorous data cleaning techniques to eliminate noisy data (e.g., irrelevant and duplicates) and ensure the uniqueness of mined data through a specific ID filtering process.
@@
Reliability: The threats to the replicability of our study primarily stem from the process of data collection, cleaning, and labeling. To alleviate these threats, we have provided a detailed description of our study design in Section 3, outlining specific steps involved in data curation. Specifically, we conducted a pilot data labeling exercise to ensure a consistent understanding of labeling criteria among authors. Any discrepancies and disagreements were resolved through discussions. Furthermore, to promote transparency and facilitate reproducibility, we have made the datasets publicly available online (Li et al., 2025). This enables researchers to independently replicate our findings and conduct further research based on our shared data.
@@ 7. Conclusions and Future Work
We provide our collected, extracted, and processed datasets online at (Li et al., 2025).
@@ Acknowledgements.
This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant Nos. 62402348 and 62172311, and the Major Science and Technology Project of Hubei Province under Grant No. 2024BAA008.

### v4

@@ 6. Threats to Validity
In this section, we discuss the potential threats to the validity of our study and the adopted measures to mitigate these threats according to the guidelines of Wohlin et al. (Wohlin et al., 2012). Internal validity is not considered in this work, as no causal relationships between variables are addressed.
Construct validity: Potential threats to this study lie in the mined ChatGPT links and their content that developers shared on GitHub. Not all shared ChatGPT links reflect meaningful interactions; some may be noise, off-topic, or duplicates. For example, many shared ChatGPT links are not relevant to software development, or certain development-related ChatGPT links are duplicated from different sources (e.g., one link occurs in both Code and Issues). To mitigate these threats, we employed rigorous data cleaning techniques to eliminate noisy data (e.g., irrelevant content and duplicates) and ensure the uniqueness of the mined data through a specific ID filtering process.
@@
Reliability: The threats to the replicability of our study primarily stem from the process of data collection, cleaning, and labeling. To alleviate these threats, we have provided a detailed description of our study design in Section 3, outlining the specific steps involved in data curation. Specifically, we conducted a pilot data labeling exercise to ensure a consistent understanding of the labeling criteria among the authors. Any discrepancies and disagreements were resolved through discussions. Furthermore, to promote transparency and facilitate reproducibility, we have made the datasets publicly available online (Li et al., 2025). This enables researchers to independently replicate our findings and conduct further research based on our shared data.
@@ 7. Conclusions and Future Work
We have provided our collected, extracted, and processed datasets of this study online at (Li et al., 2025).
@@ Acknowledgements.
This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant Nos. 92582203 and 62402348, and the Major Science and Technology Project of Hubei Province under Grant No. 2024BAA008. The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.